# GRAND for Gaussian Intersymbol Interference Channels

### 09 Mar 2026 v1

@@ III-B General Form of GRAND-ISI
${}\hat{m}=\arg\min\limits_{m=1,\cdots,M}\quad\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\gamma(\underline{y},\mathcal{I}_{t}).$ (7)
@@
  • When $\gamma(\underline{y},\mathcal{I}_{t})=r(\mathcal{I}_{t})$, where $r(\mathcal{I}_{t})$ is the rank of $\text{Rel}(\mathcal{I}_{t})$ among the sequence reliability values of all error bursts in IND, from $1$ (the smallest) to $\frac{N(N+1)}{2}$ (the largest), (7) is called ORBGRAND-ISI as an extension of ORBGRAND [26] for ISI channels.
  • When $\gamma(\underline{y},\mathcal{I}_{t})=\Psi^{-1}\left(\frac{r(\mathcal{I}_{t})}{N(N+1)/2+1}\right)$, where $\Psi(\cdot)$ is the CDF of the sequence reliability of all error bursts, (7) is CDF-ORBGRAND-ISI. The idea underlying CDF-ORBGRAND can be traced back to [26]. Moreover, [31] establishes that, over memoryless binary-input channels, CDF-ORBGRAND is exactly capacity-achieving.
@@ III-D SGRAND-ISI
If the hard detection sequence $\underline{x}^{*}$ is a codeword, SGRAND-ISI immediately outputs it as the decoding result. Otherwise, it needs to compute the sequence reliability values for all the $\frac{N(N+1)}{2}$ error bursts in IND. Facilitated by lemmas presented in the previous subsection, this computation can be significantly simplified as follows:
@@
: Compute the sequence reliability for all error bursts of size $1$;
@@
: Iteratively compute the sequence reliability for longer error bursts using Lemmas 2 and 3.
After obtaining the sequence reliability for all error bursts, we need to perform SGRAND-ISI according to (7). Candidate EPs can be generated by the heap algorithm in [25] or the parallel algorithm in [32]. In each iteration, such algorithms simply choose several error bursts from IND to produce a candidate EP. However, in light of Definition 1, for such chosen error bursts to assemble a valid EP, they should not overlap, and any two of them should be separated by at least one position. Any candidate EP not satisfying these criteria is expurgated, and only the surviving ones lead to valid EPs to be tested by parity-check matrix verification.
@@
We obtain $\underline{x}^{\star}=0001$ and the sequence reliabilities are:
@@
Based on Lemmas 2 and 3, the sequence reliability of all error bursts can be evaluated,
@@
With these, candidate EPs can be sequentially generated. Each candidate EP is represented as a length-$\frac{N(N+1)}{2}$ binary vector whose positions correspond to the elements in IND, and only those chosen error bursts have their corresponding positions marked as one. As discussed, an additional check is invoked to expurgate those invalid candidate EPs, i.e., those assembled by overlapping or consecutive error bursts. For example, the candidate EP $1001000000$ consists of error bursts $\left\{1\right\}$ and $\left\{1,2\right\}$, but since these two error bursts overlap, this candidate EP is invalid and immediately expurgated; the candidate EP $0001100000$ consists of error bursts $\left\{1,2\right\}$ and $\left\{4\right\}$, which satisfy the conditions in Definition 1, and hence this corresponds to a valid EP $1101$, which is then used for subsequent parity-check matrix verification.
@@
6:  $\mathcal{Q}_{1},\cdots,\mathcal{Q}_{\lvert\mathcal{I}(\underline{v})\rvert}\leftarrow(\underline{v},\text{IND})$
7:  $\text{t}\leftarrow$ check validity for $\underline{v}$
8:  if $t=1$ then
9:   $p=p+1$
10:   $\underline{e}\leftarrow\underline{0},\underline{e}(\bigcup_{i=1}^{\lvert\mathcal{I}(\underline{v})\rvert}\mathcal{Q}_{i})=1$
11:   if $H\cdot(\underline{x}^{*}\oplus\underline{e})^{\text{T}}=\underline{0}$ then
12:    return $\underline{\hat{x}}=\underline{x}^{*}\oplus\underline{e}$
13:   end if
14:   if $p=Q$ then
15:    return ABANDON
@@
17:  end if
18:  if $\underline{v}=\underline{0}$ then
19:   $i_{*}=0$
20:  else
21:   $i_{*}=\max\left\{i:v_{r_{i}}\neq 0\right\}$
22:  end if
23:  if $i_{*}<\operatorname{length}(\underline{r})$ then
24:   $v_{r_{(i_{*}+1)}}=1,\mathcal{S}=\mathcal{S}\cup\left\{\underline{v}\right\}$
25:   if $i_{*}>0$ then
26:    $v_{r_{i_{*}}}=0,\mathcal{S}=\mathcal{S}\cup\left\{\underline{v}\right\}$
@@
28:  end if
29:end while
30:return ABANDON
@@ III-E ORBGRAND-ISI
While SGRAND-ISI has the optimal performance as shown in Theorem 1, it requires the exact values of sequence reliability to generate EPs, hindering efficient hardware implementation. Leveraging the idea of ORBGRAND [26], we can analogously devise the ORBGRAND-ISI algorithm, with $\gamma(\underline{y},\mathcal{I}_{t})=r(\mathcal{I}_{t})$, the rank of $\text{Rel}(\mathcal{I}_{t})$ among the sequence reliability values of all error bursts in IND, from $1$ (the smallest) to $\frac{N(N+1)}{2}$ (the largest), to facilitate hardware implementation.
@@ IV-A Properties of Sequence Reliability
In the case of first-order ISI channels, the properties of sequence reliability as described by the lemmas in Section III-C enable iterative computation of sequence reliability of error bursts. For higher-order ISI channels, however, the properties are less useful than those in the first-order case, and it may sometimes be more convenient to compute sequence reliability of error bursts directly from Definition 2.
@@ IV-B Approximation Strategy
In second-order ISI channels, there are (N2)(N-2) partially-decomposable error bursts of size 2, and (N4)+2(N3)(N-4)+2(N-3) partially-decomposable error bursts of size 3 (assuming $N\geq 4$). For reducing the computational complexity, partially-decomposable error bursts of sizes larger than $3$ are not considered. After computing the sequence reliability values for non-decomposable and partially-decomposable error bursts, the subsequent operations of decoding are essentially the same as those in first-order ISI channels.
@@ V-A First-order ISI Channels
Moreover, we present the average number of queries for candidate EPs and the average number of queries for valid EPs in Fig. 3 and Fig. 4. Among all algorithms, SGRAND-ISI incurs the fewest queries, and CDF-ORBGRAND-ISI the second fewest. Although SGRAND-ISI requires fewer queries than CDF-ORBGRAND-ISI, it depends on exact values of sequence reliability to dynamically generate EPs, which in turn leads to the highest computational complexity in implementation.
@@ V-C Complexity Analysis
In the first step, the receiver employs the Viterbi algorithm to obtain the hard detection sequence, with a complexity of $\mathcal{O}(N\cdot 2^{L})$, where $N$ is code length and $L$ is ISI channel order. If the hard detection sequence is already a codeword, it is immediately declared as decoding result; otherwise, ORBGRAND-ISI proceeds to calculate and sort the sequence reliability for error bursts. The complexity of this second step is determined by the total number of sequence reliability calculations required, denoted as $\mathrm{tot}$, which can be calculated as follows:
@@
For ORBGRAND-AI, as described in [7], its procedure can also be divided into the same three steps. Therefore, we can compare its complexity with ours. Taking $\mathrm{E_{b}/N_{0}}=4$dB in Fig. 2a as an example, the results are summarized in Table II, where Step 1 refers to the complexity of the Viterbi algorithm (assuming $L\leq\text{block length}$), Step 2 denotes the number of sequence reliability values to be computed, and Step 3 reports the average number of queries (each entry is given in the form of candidate/valid, corresponding to the average numbers of queries for candidate EPs and valid EPs, respectively). We observe that ORBGRAND-ISI, 2line-CDF-ORBGRAND-ISI, and CDF-ORBGRAND-ISI incur similar computational complexity. Compared to ORBGRAND-AI12, the number of computed sequence reliability values is reduced by about a factor of five, and the corresponding number of queries is also reduced accordingly.
@@ VII-A Proof of Theorem 1
Denote $\mathcal{B}_{t}(m)=\mathcal{I}_{t}\cup\left\{\text{R}_{1}(\mathcal{I}_{t})\right\}$ for $t=1,\cdots,\zeta(\mathcal{I}(\underline{e}(m)))$ and $\mathcal{B}(m)=\mathcal{B}_{1}(m)\cup\mathcal{B}_{2}(m)\cdots\cup\mathcal{B}_{\zeta(\mathcal{I}(\underline{e}(m)))}(m)$. Based on the defintion of (5), we have
$\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\text{Rel}(\mathcal{I}_{t})$
$=$ $\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\sum_{i=\min(\mathcal{I}_{t})}^{\text{R}_{1}(\mathcal{I}_{t})}\ln\frac{P_{\mathsf{Y}_{i}|\underline{\mathsf{X}}_{i-1:i}}(y_{i}|\underline{x}^{*}_{i-1:i})}{P_{\mathsf{Y}_{i}|\underline{\mathsf{X}}_{i-1:i}}(y_{i}|[f_{\mathcal{I}_{t}}(\underline{x}^{*})]_{i-1:i})}$
$=$ $\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\sum_{i=\min(\mathcal{I}_{t})}^{\text{R}_{1}(\mathcal{I}_{t})}\ln\frac{P_{\mathsf{Y}_{i}|\underline{\mathsf{X}}_{i-1:i}}(y_{i}|\underline{x}^{*}_{i-1:i})}{P_{\mathsf{Y}_{i}|\underline{\mathsf{X}}_{i-1:i}}(y_{i}|\underline{x}(m)_{i-1:i})}$
@@
$\hat{m}$= $\arg\min\limits_{m=1,\cdots,M}\quad\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\gamma(\underline{y},\mathcal{I}_{t})$

### 10 Mar 2026 v2

@@ III-B General Form of GRAND-ISI
${}\hat{m}=\arg\min\limits_{m=1,\cdots,M}\quad\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\gamma\big(\underline{y},\mathcal{I}_{t}(\underline{e}(m))\big).$ (7)
@@
  • When $\gamma(\underline{y},\mathcal{I}_{t})=r(\mathcal{I}_{t})$, where $r(\mathcal{I}_{t})$ is the rank of $\text{Rel}(\mathcal{I}_{t})$ among the sequence reliabilities of all error bursts in IND, from $1$ (the smallest) to $\frac{N(N+1)}{2}$ (the largest), (7) is called ORBGRAND-ISI as an extension of ORBGRAND [26] for ISI channels.
  • When $\gamma(\underline{y},\mathcal{I}_{t})=\Psi^{-1}\left(\frac{r(\mathcal{I}_{t})}{N(N+1)/2+1}\right)$, where $\Psi(\cdot)$ is the CDF of the sequence reliabilities of all error bursts, (7) is CDF-ORBGRAND-ISI. The idea underlying CDF-ORBGRAND can be traced back to [26]. Moreover, [31] establishes that, over memoryless binary-input channels, CDF-ORBGRAND is exactly capacity-achieving.
@@ III-D SGRAND-ISI
If the hard detection sequence $\underline{x}^{*}$ is a codeword, SGRAND-ISI immediately outputs it as the decoding result. Otherwise, it needs to compute the sequence reliabilities for all the $\frac{N(N+1)}{2}$ error bursts in IND. Facilitated by lemmas presented in the previous subsection, this computation can be significantly simplified as follows:
@@
: Compute the sequence reliabilities for all error bursts of size $1$;
@@
: Iteratively compute the sequence reliabilities for longer error bursts using Lemmas 2 and 3.
After obtaining the sequence reliabilities for all error bursts, we need to perform SGRAND-ISI according to (7). Candidate EPs can be generated by the heap algorithm in [25] or the parallel algorithm in [32]. In each iteration, such algorithms simply choose several error bursts from IND to produce a candidate EP. However, in light of Definition 1, for such chosen error bursts to assemble a valid EP, they should not overlap, and any two of them should be separated by at least one position. Any candidate EP not satisfying these criteria is expurgated, and only the surviving ones lead to valid EPs to be tested by parity-check matrix verification.
@@
We obtain $\underline{x}^{\star}=0001$ and the sequence reliabilities for all error bursts of size $1$ are:
@@
Based on Lemmas 2 and 3, the sequence reliabilities of all error bursts can be evaluated,
@@
[…]
With these, candidate EPs can be sequentially generated. Each candidate EP is represented as a length-$\frac{N(N+1)}{2}$ binary vector whose positions correspond to the elements in IND, and only those chosen error bursts have their corresponding positions marked as one. As discussed, an additional check is invoked to expurgate those invalid candidate EPs, i.e., those assembled by overlapping or consecutive error bursts. For example, the candidate EP $1001000000$ consists of error bursts $\left\{3\right\}$ and $\left\{2,3\right\}$, but since these two error bursts overlap, this candidate EP is invalid and immediately expurgated; the candidate EP $0000011000$ consists of error bursts $\left\{1,2\right\}$ and $\left\{4\right\}$, which satisfy the conditions in Definition 1, and hence this corresponds to a valid EP $1101$, which is then used for subsequent parity-check matrix verification.
@@
6:  $\text{t}\leftarrow$ check validity for $\underline{v}$
7:  if $t=1$ then
8:   $p=p+1$
9:   $\underline{e}\leftarrow(\underline{v},\text{IND})$
10:   if $H\cdot(\underline{x}^{*}\oplus\underline{e})^{\text{T}}=\underline{0}$ then
11:    return $\underline{\hat{x}}=\underline{x}^{*}\oplus\underline{e}$
12:   end if
13:   if $p=Q$ then
14:    return ABANDON
15:   end if
@@
17:  if $\underline{v}=\underline{0}$ then
18:   $i_{*}=0$
19:  else
20:   $i_{*}=\max\left\{i:v_{r_{i}}\neq 0\right\}$
21:  end if
22:  if $i_{*}<\operatorname{length}(\underline{r})$ then
23:   $v_{r_{(i_{*}+1)}}=1,\mathcal{S}=\mathcal{S}\cup\left\{\underline{v}\right\}$
24:   if $i_{*}>0$ then
25:    $v_{r_{i_{*}}}=0,\mathcal{S}=\mathcal{S}\cup\left\{\underline{v}\right\}$
26:   end if
@@
28:end while
29:return ABANDON
@@ III-E ORBGRAND-ISI
While SGRAND-ISI has the optimal performance as shown in Theorem 1, it requires the exact value of sequence reliability to generate EPs, hindering efficient hardware implementation. Leveraging the idea of ORBGRAND [26], we can analogously devise the ORBGRAND-ISI algorithm, with $\gamma(\underline{y},\mathcal{I}_{t})=r(\mathcal{I}_{t})$, the rank of $\text{Rel}(\mathcal{I}_{t})$ among the sequence reliabilities of all error bursts in IND, from $1$ (the smallest) to $\frac{N(N+1)}{2}$ (the largest), to facilitate hardware implementation.
@@ IV-A Properties of Sequence Reliability
In the case of first-order ISI channels, the properties of sequence reliability as described by the lemmas in Section III-C enable iterative computation of sequence reliabilities of error bursts. For higher-order ISI channels, however, the properties are less useful than those in the first-order case, and it may sometimes be more convenient to compute sequence reliabilities of error bursts directly from Definition 2.
@@ IV-B Approximation Strategy
In second-order ISI channels, there are (N2)(N-2) partially-decomposable error bursts of size 2, and (N4)+2(N3)(N-4)+2(N-3) partially-decomposable error bursts of size 3 (assuming $N\geq 4$). For reducing the computational complexity, partially-decomposable error bursts of sizes larger than $3$ are not considered. After computing the sequence reliabilities for non-decomposable and partially-decomposable error bursts, the subsequent operations of decoding are essentially the same as those in first-order ISI channels.
@@ V-A First-order ISI Channels
Moreover, we present the average number of queries for candidate EPs and the average number of queries for valid EPs in Fig. 3 and Fig. 4. Among all algorithms, SGRAND-ISI incurs the fewest queries, and CDF-ORBGRAND-ISI the second fewest. Although SGRAND-ISI requires fewer queries than CDF-ORBGRAND-ISI, it depends on exact values of sequence reliabilities to dynamically generate EPs, which in turn leads to the highest computational complexity in implementation.
@@ V-C Complexity Analysis
In the first step, the receiver employs the Viterbi algorithm to obtain the hard detection sequence, with a complexity of $\mathcal{O}(N\cdot 2^{L})$, where $N$ is code length and $L$ is ISI channel order. If the hard detection sequence is already a codeword, it is immediately declared as decoding result; otherwise, ORBGRAND-ISI proceeds to calculate and sort the sequence reliabilities for error bursts. The complexity of this second step is determined by the total number of sequence reliability calculations required, denoted as $\mathrm{tot}$, which can be calculated as follows:
@@
For ORBGRAND-AI, as described in [7], its procedure can also be divided into the same three steps. Therefore, we can compare its complexity with ours. Taking $\mathrm{E_{b}/N_{0}}=4$dB in Fig. 2a as an example, the results are summarized in Table II, where Step 1 refers to the complexity of the Viterbi algorithm (assuming $L\leq\text{block length}$), Step 2 denotes the number of sequence reliabilities to be computed (or reliabilities in ORBGRAND-AI), and Step 3 reports the average number of queries (each entry is given in the form of candidate/valid, corresponding to the average numbers of queries for candidate EPs and valid EPs, respectively). We observe that ORBGRAND-ISI, 2line-CDF-ORBGRAND-ISI, and CDF-ORBGRAND-ISI incur similar computational complexity. Compared to ORBGRAND-AI12, the number of computed sequence reliabilities is reduced by about a factor of five, and the corresponding number of queries is also reduced accordingly.
@@ VII-A Proof of Theorem 1
Denote $\mathcal{B}_{t}(m)=\mathcal{I}_{t}(\underline{e}(m))\cup\left\{\text{R}_{1}\big(\mathcal{I}_{t}(\underline{e}(m))\big)\right\}$ for $t=1,\cdots,\zeta(\mathcal{I}(\underline{e}(m)))$ and $\mathcal{B}(m)=\mathcal{B}_{1}(m)\cup\mathcal{B}_{2}(m)\cdots\cup\mathcal{B}_{\zeta(\mathcal{I}(\underline{e}(m)))}(m)$. Based on the defintion of (5), we have
$\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\text{Rel}(\mathcal{I}_{t}(\underline{e}(m)))$
$=$ $\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\sum_{i=\min(\mathcal{I}_{t}(\underline{e}(m)))}^{\text{R}_{1}(\mathcal{I}_{t}(\underline{e}(m)))}\ln\frac{P_{\mathsf{Y}_{i}|\underline{\mathsf{X}}_{i-1:i}}(y_{i}|\underline{x}^{*}_{i-1:i})}{P_{\mathsf{Y}_{i}|\underline{\mathsf{X}}_{i-1:i}}(y_{i}|[f_{\mathcal{I}_{t}}(\underline{x}^{*})]_{i-1:i})}$
$=$ $\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\sum_{i=\min(\mathcal{I}_{t}(\underline{e}(m)))}^{\text{R}_{1}(\mathcal{I}_{t}(\underline{e}(m)))}\ln\frac{P_{\mathsf{Y}_{i}|\underline{\mathsf{X}}_{i-1:i}}(y_{i}|\underline{x}^{*}_{i-1:i})}{P_{\mathsf{Y}_{i}|\underline{\mathsf{X}}_{i-1:i}}(y_{i}|\underline{x}(m)_{i-1:i})}$
@@
$\hat{m}$= $\arg\min\limits_{m=1,\cdots,M}\quad\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\gamma\big(\underline{y},\mathcal{I}_{t}(\underline{e}(m))\big)$
$=$ $\arg\min\limits_{m=1,\cdots,M}\quad\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\text{Rel}(\mathcal{I}_{t}(\underline{e}(m)))$